### v1

@@ 5 Discussion and conclusion
We use the near-optimal feasible space of an energy system model to generate 56,050 alternative energy system designs for Longyearbyen, Svalbard. The alternative system configurations that we generated were then used in an interactive interface, with which 126 local inhabitants explored the implications of investing in different energy system components. This allowed the participants to explore the trade-offs between different options. Compared to previous interactive modelling tools, such as the Riskmeter [17, 20], utilising near-optimal model results facilitates a deeper model integration, where the information presented to the participants includes the system interactions of different components and the intermittency of variable renewable electricity generation. One critical feature is that the decision space for a user is limited to only feasible system configurations, which meet all the modelling constraints. This prevents users from selecting a combination of system components that, for example, will fail to supply energy at all time.
As illustrated in figure 4 and from the stated priorities of the participants, it becomes apparent that people care for more than only costs when evaluating future energy systems. Most participants in our study state that it is a balance between three or more priorities, hinting at the complexity of several simultaneous trade-offs when making decisions. This is an important consideration for the interaction of energy system modelling and policy-making, which has traditionally focused on cost-optimal solutions, potentially overlooking near-optimal solutions which may be socially more acceptable for other reasons. Presenting stakeholders with near-optimal modelling results allows for interactions and interfaces that give the users considerable diversity of options and flexibility when exploring trade-offs and identifying systems that align with their priorities. At the same time, it enables stakeholder interactions that build on realistic alternatives and system knowledge generated from energy system models.
The five dimensions (slider options) that the participants were able to directly influence were meant to represent the most relevant technologies, following the local energy transition plan [40]. Although this constrained the participants to select system configurations that were both technologically and politically feasible, participants mentioned technologies that did not meet these criteria, both during the process and in the post-study feedback. Mentions included a return to the recently discontinued coal-fired thermal power plant, nuclear power and geothermal energy. Longyearbyen was founded as a coal mining settlement and coal has played an important part in forming the identity of the town as well as legitimising Norwegian presence in Svalbard [41]. Despite this, coal power was excluded from the model and interface based on the political decisions that the energy supply in Longyearbyen should support and be in line with Norway’s climate targets for 2030 and 2050 [42]. Nuclear is not considered techno-economically feasible, as both conventional nuclear and small modular reactors would be over-dimensioned in Longyearbyen. The fact that participants are interested in technologies and solutions that are considered infeasible hints at the usefulness of these kind of exercises, both to understand why opinions are formed and ensure that the opinions are informed.
Furthermore, some participants also stated that demand/consumption reduction was an important measure that should be considered. Of course, reducing the energy consumption would reduce the need for new infrastructure development and lower the negative implications for the system. This could be an interesting dimension to include in such an interactive interface, but the way it is presented would need to be carefully considered. Simply reducing the energy needs may seem very desirable from a system perspective, but the implications on both society and individuals would need to be illustrated in an effective manner.
Although the methodology is adaptable to other contexts and environments, some of the results need to be seen as particular to the context of our study. The energy system in Longyearbyen is relatively small and more familiar than for the average city, and the future energy system has been part of the public discourse for many years, leading to an expectation that the participants are relatively well-informed on the different alternatives and implications. The remoteness and low temperatures in this Arctic settlement together with geopolitical tension in the area makes it susceptible to various types of interruptions. This likely leads to a higher valuation of vulnerability than other locations.
@@
Whereas these results are from a particular case study in Longyearbyen, Svalbard, the approach and methodology can be extended to other circumstances. The model and particularly the interface is customisable. However, modelling and studying local communities that are not isolated like Longyearbyen present some challenges that need to be addressed, such as the option of importing electricity from neighbouring areas. Future applications should also consider engaging with stakeholders in different ways. It may be beneficial with fewer participants who spend more time with the interface, in order to better understand the processes and underlying methods. Although several of our participants stated that the process was insightful and that they enjoyed the interface, there are also mentions of that it was difficult to understand exactly what was going on, as well as suggestions of making the interface available on the local government’s website, for people to use in their own time. Whereas we achieved a good number of participants and good diversity in terms of demographics, approaching them in public spaces limited how much time they had available to get acquainted with the interface and how it worked. This became evident from the fact that no one, to our knowledge, utilised or read the information sheet that we prepared and made available to the participants. Instead participants preferred to directly enquire for specific explanations instead of a full overview of the modelling assumptions. Future work should also look to include stakeholders throughout the processes, and not only at one stage, as with our particular application. A greater degree of participation may be achieved by, for example, co-designing the dimensions (decision-variables) and relevant metrics together with stakeholders.
The interest for near-optimality in energy systems modelling has grown considerably in the last few years [24, 28] and new advancements are made, which provide opportunities for further fine-tuning of the methodology presented here. For example, Lau et al. [43] introduced a novel MGA post-processing method, Modelling to Generate Continuous Alternatives, which allows for detailed exploration of the near-optimal feasible space, with a dimensionality reduction method that could be run “live” as a response to stakeholder input, potentially drastically reducing the need for pre-computations and assumptions made by the modellers. Lombardi and Pfenninger [29] propose a Human-in-the-loop (HITL) extension to conventional MGA approaches, which generates a secondary collection of near-optimal alternatives closely aligned with initially selected preferred solutions. These, and other, improvements to MGA and near-optimality in energy system models may contribute to further enhancing its usefulness for stakeholder interactions, elicitation of informed stakeholder preferences and as an information/learning tool.
@@
The maximum slack level applied in any near-optimal workflow will naturally have a large impact on the energy system configurations possible. Whereas many other studies commonly apply values of up to 10-15% [44, 45], we purposefully include a much higher slack level (125%) to broaden the diversity of options presented to participants, and instead let them decide the slack level of their system. The stated willingness to pay from the post-study questionnaire has a median of 15%, similar to both previous studies and within the range of how close to cost-optimal previous transitions have been (9-23%) [46]. On the other hand, the revealed willingness to pay from the submitted energy system designs has a median of 91%. This indicate that even though participants have a preconception of how much they are willing to deviate from cost-optimality, the energy systems that they seek may require compromises with respect to its additional cost. Further research into this topic may reveal realistic slack values for future near-optimal modelling.
@@
For the energy transition to succeed, social and political feasibility is of utmost importance. Contested technologies and infrastructure risk slowing down processes and limit political will. Supporting legitimacy in decisions and processes, by allowing for people to make their voices heard as such facilitate energy transitions and improves our chance of collectively limiting climate change and staying within the planetary boundaries. We have, with this work, demonstrated a methodology for combining near-optimality with stakeholder interactions that show promise for illustrating and navigating difficult trade-offs in energy systems planning. Participants who have interacted with the methodology have clearly shown that there is more to consider than only cost-optimality, and that it is often a mix of several important decisions which cannot all be met. Engaging with a continuum of feasible system designs via visualisation tools, such as our interactive interface, can help demonstrate the complexity in planning decisions, and also create meaning and a positive feeling around the challenges and dilemmas of energy transitions.
@@ 6 Methods
Data collection
@@
The 126 participants represent Longyearbyen fairly well in several categories. There is a balance between male and female participants, different age groups and both those with an without Norwegian citizenship. Most of the participants had not stayed in Longyearbyen for a very long time, which is common in the very transient population of Longyearbyen. A few participants had been there for 10, 20 or even 30 years.
@@
Data and code availability The source code for the interactive web interface is available on GitHub (https://github.com/koen-vg/near-opt-interface/) and the code for generating the near-optimal solutions used in the interactive web interface will be made available in the next version of this document.

### v2

@@ 5 Discussion and conclusion
We use the near-optimal feasible space of an energy system model to generate 56,050 alternative energy system designs for Longyearbyen, Svalbard. These system configurations were presented to 126 local citizens who explored different investments in energy system components through an interactive interface. It allowed the participants to explore trade-offs between and implications of different options. Compared to previous interactive modelling tools [18, 21], the deeper integration through near-optimal modelling incorporated system interactions of different components as well as intermittency of renewables. Most critically, this approach limits the decision space to feasible system configurations. This prevents users from selecting systems that fail to reliably supply energy, which is of utmost importance in vulnerable communities like Longyearbyen.
Stated and revealed preferences in the modelling (also figure 4) demonstrate that participants do not only care about the cost of their energy (system). Most participants state to balance at least three priorities, hinting at the complexity of several simultaneous trade-offs when making decisions. The common, narrow focus on cost optimality in energy modelling and policymaking is therefore potentially overlooking near-optimal solutions with higher social acceptance. The interface we present gives stakeholders the opportunity to interact with options and trade-offs, as well as to explore their priorities and how to achieve them. At the same time, it adds the system knowledge and techno-economic value of energy system modelling.
The five dimensions/sliders that the participants were adjusting were meant to represent the most relevant technologies, following the local energy transition plan [41]. Although the modelling reflected political and technological realities, both during the process and in the post-study feedback, participants mentioned technologies such as coal, nuclear and geothermal energy that did not meet these criteria. Mentions of returning to the recently discontinued coal-fired thermal power plant can be tied to Longyearbyen’s history as a coal mining settlement which formed local identity and legitimised Norwegian presence on Svalbard [42]. The interest in technologies that are considered infeasible (see Methods as to why) highlights the importance of participatory studies as they reveal opinions, and help to understand why they are formed and keep society informed. Furthermore, some participants stated a preference for reduced demand / consumption, which could indeed reduce trade-offs and repercussions of new energy infrastructure developments. This would be an interesting dimension to include in such an interface, but how it is presented needs to be carefully considered: from a system perspective it may appear desirable, but societal and individual impacts would need to be illustrated effectively.
@@
The interface and model are highly customisable and many insights are transferable to other circumstances; some challenges remain and certain results may be specific to Longyearbyen. Participants in Longyearbyen are expected to be relatively well-informed: not only is the energy system small and familiar, its transition has also been publicly discussed for many years. The remoteness and bitterly cold climate together with geopolitical tensions expose risks to various types of interruptions, and a greater degree of vulnerability than elsewhere. Besides, most local communities might not directly experience all impacts of their decisions, e.g. if they are connected to neighbouring areas and can import energy. Moreover, future applications should consider engaging with participants who spend more time with the interface to enhance the understanding of the processes and methods. Whereas we achieved a good number of participants and good diversity in terms of demographics, through the interactions in public spaces, this limited the time available for participants to get acquainted with the interface and its workings. A few users were concerned about the complexity of the underlying modelling, whereas others stated that it was insightful and enjoyable and suggested making the interface available on the community council’s website. It should be noted, however, that no participant, to our knowledge, utilised or read the information sheet that we prepared and laid out. Instead of a full overview of the modelling assumptions, participants preferred direct enquiries for specific explanations. Future work should also look to include stakeholders throughout the processes, and not only at a single stage, as with our particular application. A greater degree of participation may be achieved by, for example, co-designing the dimensions (decision-variables) and relevant metrics together with stakeholders.
@@
Furthermore, the slack or admissible cost increase in near-optimal workflows impacts which energy system configurations are possible. In contrast to common values of up to 10-15% [43, 44], we purposefully increase slack levels (125%) to broaden the diversity of options, and instead let participants decide the slack level of their system. The stated willingness-to-pay from the post-study questionnaire has a median of 15%, similar to both previous studies and within the range of how close to cost-optimal previous transitions have been (9-23%) [45]. On the other hand, the revealed willingness-to-pay from the submitted energy system designs has a median of 91%. This indicates tensions between preconceptions on “acceptable” deviations from cost-optimality and the realisation of their own preferences. Further research into this topic may reveal realistic slack values for future near-optimal modelling.
Growing interest [25, 34] and recent advances of near-optimal methods in energy system modelling provide opportunities for further fine-tuning of the presented methodology. For example, an MGA post-processing by Lau et al. [46] can respond “live” to stakeholder input when exploring the near-optimal space, potentially reducing pre-computations as well as modelling assumptions. Lombardi and Pfenninger [35] propose a Human-in-the-loop (HITL) extension to conventional MGA approaches, which generates a secondary collection of near-optimal alternatives closely aligned with initially selected preferred solutions. These, and other improvements to MGA and near-optimality in energy system models may contribute to further enhancing its usefulness for stakeholder interactions, elicitation of informed stakeholder preferences and as a tool for learning.
@@
For the energy transition to succeed, social and political feasibility is of utmost importance. Contested technologies and infrastructure risk slowing down processes and limit political will. Supporting legitimacy in decisions and processes, and making people’s voices heard in energy transitions improves our chance of collectively limiting climate change and staying within the planetary boundaries. We have, with this work, demonstrated a methodology for combining near-optimality with stakeholder interactions that shows promise for illustrating and navigating difficult trade-offs in energy systems planning. Participants who have interacted with the methodology have clearly shown that there is more to consider than only cost-optimality, and that all preferences cannot always be met. Engaging with a continuum of feasible system designs via visualisation tools, such as our interactive interface, can help demonstrate the complexity in planning decisions, and also create meaning and a positive feeling around the challenges and dilemmas of energy transitions.
@@ 6 Methods
Coal power was excluded from the model and interface based on the political decisions that the energy supply in Longyearbyen should support and be in line with Norway’s climate targets for 2030 and 2050 [47]. Nuclear is not considered techno-economically feasible, as both conventional nuclear and small modular reactors would be over-dimensioned in Longyearbyen.
@@
Data collection and field work
@@
The 126 participants represent Longyearbyen fairly well in several categories. There is a balance between male and female participants, different age groups and both those with and without Norwegian citizenship. Most of the participants had not stayed in Longyearbyen for a very long time, which is common in the very transient population of Longyearbyen. A few participants had been there for 10, 20 or even 30 years.
@@
Author contributions All authors conceived the idea and designed the study. O.V., A.G. and K.vG. did the field work and collected the data. A.G. lead the implementation of the near-optimal modelling. K.vG. lead the development of the interactive interface. O.V. lead the writing of the paper, with considerable input from the other authors.
[…]
Data and code availability The source code for the interactive web interface and the code for generating the near-optimal solutions used in the interactive web interface is openly available on GitHub: interactive interface: https://github.com/koen-vg/near-opt-interface/ and near-optimal workflow: https://github.com/aleks-g/lyb.
The field data that support part of the findings of this study are not available owing to the confidentiality of participants’ identities. Longyearbyen is a small settlement, and providing individual data points may risk the identification of particular individuals, even if the data is anonymised. Aggregated data can be made available upon request.